OpenStudy (anonymous):

A fire hose has an opening whose radius is 10 cm. The hose discharges water at a velocity of 20 m/s. How much force is required by the fireman to hold the nozzle stationary? A fire hose ejects a stream of water with a speed of 20 m/s, as above. At what angle to the horizontal should the fire hose be directed if it is to remove the civil liberties of protesters 20 m from where the hose is being held?

OpenStudy (anonymous):

@KingGeorge @robtobey @Yuba hello all, please help me with this problem. thanks

OpenStudy (anonymous):

anyone ? :s please help me @erickkoro @gyanu @hoblos @linda3 @ybarrap @ZeHanz

OpenStudy (anonymous):

@beccaboo333 thanks

OpenStudy (dumbsearch2):

Required force, F = 2*pi*((0.10)^2}*20*1000*9.81 N [pi*r^2*v*g*(density of water, in kg/m^3) or F = 135604 N angle theta is given by {(20^2)/9.81}sin 2*(theta) = 20 or sin (2*theta) = 9.81/20 or 2*theta = 0.4905 or 2*theta = 29.37 degree or theta = 14.7 degree

OpenStudy (anonymous):

Thanks a lot mate.